OpenStudy (anonymous):

Concrete can be purchased by the cubic yard. How much will it cost to pour a slab 18 feet by 18 feet by 6 inches for a patio if the concrete costs $42.00 per cubic yard? (1 point)

OpenStudy (campbell_st):

easiest think to do is convert all the measurements to yards... 18ft = 6 yards 6 inch = 1/6 yard ( 6 inches = 1/2 ft.... so 1/6 yard) so the volume is V = 6 x 6 x 1/6 then calculate the cost hope this helps